# Response of superfluid fermions at finite temperature

### v3

@@ Abstract
A consistent finite-temperature microscopic theory for the response of strongly coupled superfluid fermionic systems is formulated. We start from the general many-body Hamiltonian with the vacuum (bare) two-fermion interaction and derive the equation of motion (EOM) for the thermally averaged two-time two-fermion correlation function, which determines the spectrum of the system under study. The superfluidity is introduced via the Bogoliubov transformation of the fermionic field operators, and the entire formalism is carried out in the basis of Bogoliubov’s quasiparticles, keeping the complete $4\times 4$ block matrix structure of the two-fermion EOM. Static and dynamical interaction kernels of the resulting EOM are discussed. The least correlated static kernel reproduces the known finite-temperature quasiparticle random phase approximation, while the dynamical kernel was previously largely unknown in the finite-temperature superfluid theory. The major focus is then placed on the latter kernel, which is completed and advanced to a factorized form, enabling a minimal truncation of the many-body problem with keeping important effects of emergent collectivity and mapping to the quasiparticle-vibration coupling (qPVC). As in the zero-temperature and non-superfluid cases, the qPVC can be associated with a new order parameter, qPVC vertex. In the thermal superfluid theory, the latter vertex as well as the components of the dynamical kernel acquire an extended form including thermally unblocked transition amplitudes.
@@ I Introduction
Response functions can be conveniently related to correlation functions (CFs) of the time-dependent fermionic field operators and play the role of scattering amplitudes in correlated media at zero and finite temperatures. Such CFs form the common formal quantum field theory (QFT) background across the energy scales, from particle physics to quantum chemistry. Many-body systems with medium and strong coupling require non-perturbative approaches, in contrast to weakly coupled ones, which can be reasonably tackled by perturbation theory with a small parameter, typically associated with the interaction strength. A non-perturbative treatment can be organized by rearranging the many-body Hamiltonian in a way the perturbation theory can be still applied or, alternatively, by different criteria. In both cases, an adequate quantification of microscopic mechanisms of emergent collectivity and its thermal evolution plays an important role in solving the quantum many-body problem at different scales.
@@
It is recognized that the nuclear response theory capable of quantitatively addressing nuclear spectra must contain in-medium induced NN interaction which is beyond the reach of the random phase approximation (RPA) Bohm and Pines (1951), or its superfluid variant, the quasiparticle RPA (QRPA) Baranger (1960); Ring and Schuck (1980). The leading effects are attributed to the minimal coupling between the single-particle and emergent collective degrees of freedom Bohr and Mottelson (1969, 1975); Broglia and Bortignon (1976); Bortignon et al. (1977); Bertsch et al. (1983); Soloviev (1992) which cause fragmentation of the (Q)RPA states or damping of collective excitations. On the formal level, the quasiparticle-vibration coupling (qPVC) can be linked to the dynamical kernels of the equations of motion (EOM) for the CFs in nuclear medium Ring and Schuck (1980); Schuck (1976); Adachi and Schuck (1989); Danielewicz and Schuck (1994); Dukelsky et al. (1998); Schuck and Tohyama (2016); Litvinova and Schuck (2019, 2020). In the regime of the NN intermediate coupling, qPVC emerges with a new order parameter associated with the qPVC vertex, which can be used to produce a hierarchy of approximations to nucleonic propagators. The leading-order qPVC kernels can be mapped to the kernels of nuclear field theory (NFT) and related to the formalism of the quasiparticle-phonon models (QPM) Bohr and Mottelson (1969, 1975); Broglia and Bortignon (1976); Bortignon et al. (1977); Bertsch et al. (1983); Tselyaev (1989); Kamerdzhiev et al. (1997); Tselyaev (2013); Soloviev (1992) in the fully microscopic EOM framework Litvinova and Schuck (2019, 2020). The latter allowed for establishing a connection between the ab initio and effective theories and extending NFT to more complex configurations.
@@
In Ref. Litvinova and Zhang (2022), the two-fermion EOM approach was advanced to the superfluid case by working out the complete formalism in the space of the Bogoliubov quasiparticles and keeping the 2$\times$2 matrix structure of the Hartree-Fock-Bogoliubov (HFB) basis. As in the non-superfluid phase, the modification of the bare interaction in the strongly correlated medium was derived, revealing the splitting of the interaction kernel into the static and dynamical components, in principle, calculable from the underlying bare interactions. The static kernel is associated with short-range correlations, while its dynamical counterpart generates long-range correlations originating from the retardation effects of the correlated medium, and these correlations, in turn, affect the static kernel. The major theoretical step made in Ref. Litvinova and Zhang (2022) was the unification of the particle-hole ($ph$) and particle-particle ($pp$) channels in one EOM with the interaction kernel unifying the normal and pairing phonons and thus extending the applicability of the theory to open-shell nuclei which constitute the major part of the nuclear landscape.
@@
The finite-temperature variants of the response theory beyond QRPA were worked out for the $ph$ and $pp$ channels separately in Refs. Litvinova and Wibowo (2018); Wibowo and Litvinova (2019); Litvinova et al. (2020) and Litvinova and Schuck (2021), respectively. Applications to the electric dipole strength Litvinova and Wibowo (2018); Wibowo and Litvinova (2019), beta decay Litvinova et al. (2020) and electron capture rates Litvinova and Robin (2021) revealed the necessity of completion of the theory by developing a finite-temperature formalism including superfluidity in a unified way. This is done in the present work coherently. We derive and discuss the most general EOM and leading approximations to the interaction kernel. We benchmark the static kernel of the superfluid response by the finite-temperature QRPA Sommermann (1983); Ring et al. (1984) and then focus on the qPVC variants of the dynamical kernel bringing them to the form ready for numerical implementations.
@@ II.1.1 Zero temperature
with ${\bar{v}}_{1234}=v_{1234}-v_{1243}$ being the antisymmetrized matrix elements of the interaction between two fermions in free space. For a self-bound many-body system, $V$ in the analytical form or in the form of the two-body matrix elements should define the properties of the system, in principle, uniquely. Here we assume only the instantaneous character of the interaction without detailing its other properties. This assumption should be reasonable enough for the low-energy fermionic systems of interest. A superfluid system of fermions is most conveniently described in the space of Bogolyubov’s quasiparticles, which mix particle and hole states via the transformation Bogolyubov (1958); Ring and Schuck (1980)
@@
[…]
@@
$+{\tilde{F}}^{11}_{\mu\mu^{\prime}}\alpha^{\dagger}_{\mu^{\prime}}\alpha_{\mu}\Bigr{)}+\sum_{\mu}F^{00}_{\mu},$
@@
×$\times\left(\begin{array}[]{cccc}X^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] Y^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] U^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] V^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\end{array}\right)\left(\begin{array}[]{c}F^{02\ast}_{\nu\nu^{\prime}}\\[3.0pt] F^{20\ast}_{\nu\nu^{\prime}}\\[3.0pt] F^{11\ast}_{\nu\nu^{\prime}}\\[3.0pt] {\tilde{F}}^{11\ast}_{\nu\nu^{\prime}}\end{array}\right)$ (29)
@@
Thanks to the block matrix notations, the EOM for ${\hat{R}}$ can be derived keeping the $4\times 4$ structure consistently. Namely, we consider the derivative with respect to the first time argument:
@@
$[{H},A](t)=e^{i{H}t}[{H},A]e^{-i{H}t}$ (44)
for an arbitrary single-time operator (product) $A$. At this step, we evaluate explicitly the commutators with the one-quasiparticle part of the full Hamiltonian (13):
@@
$[H,{\hat{\Psi}}_{{\mu\mu^{\prime}}}]=-{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}+[V,{\hat{\Psi}}_{{\mu\mu^{\prime}}}]$ (46)
@@
$+i\langle T[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](t){\hat{\Psi}}^{\dagger}_{\nu\nu^{\prime}}(t^{\prime})\rangle,$ (55)
@@
${\hat{F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t$displaystyle-- $t^{\prime})[-i\overleftarrow{\partial_{t^{\prime}}}-{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}]=$ (59)
@@
where ${\hat{T}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ and ${\hat{T}}^{r}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t-t^{\prime})$ introduce the generalized static and dynamical (retarded, or time-dependent) parts of the two-fermion ${\hat{T}}$-matrix in the quasiparticle space:
@@
$\left[i\partial_{t}\right.$displaystyle-- $\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}\right]{\hat{R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t-t^{\prime})\left[-i\overleftarrow{\partial_{t^{\prime}}}-{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}\right]$ (62)
@@
$\left[\omega\right.$displaystyle-- $\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}\right]{\hat{R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega)\left[\omega\right.-\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}\right]$
@@
$+\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{R}}^{0}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}(\omega){\hat{T}}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}(\omega){\hat{R}}^{0}_{\delta\delta^{\prime}\nu\nu^{\prime}}(\omega)$ (65)
@@
Transformation of the $T$-matrix equation to the Dyson-Bethe-Salpeter equation (Dyson-BSE) form can be performed by introducing the interaction kernel ${\hat{K}}(\omega)={\hat{K}}^{0}+{\hat{K}}^{r}(\omega)$, irreducible with respect to ${\hat{R}}^{0}$, i.e. ${\hat{K}}$ does not contain terms partitioned by the uncorrelated propagator ${\hat{R}}^{0}$. The interaction kernel plays the role of the self-energy for the two-point two-quasiparticle correlation function:

### v4

@@ Abstract
A consistent finite-temperature microscopic theory for the response of strongly coupled superfluid fermionic systems is formulated. We start from the general many-body Hamiltonian with the vacuum (bare) two-fermion interaction and derive the equation of motion (EOM) for the thermally averaged two-time two-fermion correlation function, which determines the spectrum of the system under study. The superfluidity is introduced via the Bogoliubov transformation of the fermionic field operators, and the entire formalism is carried out in the basis of Bogoliubov’s quasiparticles, keeping the complete $4\times 4$ block matrix structure of the two-fermion EOM. Fully correlated static and dynamical interaction kernels of the resulting EOM are discussed. A special focus is then placed on the latter kernel, which is advanced to a factorized form, enabling a minimal truncation of the many-body problem while keeping important effects of emergent collectivity and mapping to the quasiparticle-vibration coupling (qPVC). As in the zero-temperature and non-superfluid cases, the qPVC can be associated with a new order parameter, qPVC vertex. In the thermal superfluid theory, the latter vertex as well as the components of the dynamical kernel acquire an extended form including thermally unblocked transition amplitudes.
@@ I Introduction
Response functions can be conveniently related to correlation functions (CFs) of the time-dependent fermionic field operators and play the role of scattering amplitudes in correlated media at zero and finite temperatures. Such CFs form the common formal quantum field theory (QFT) background across the energy scales, from particle physics to quantum chemistry. Many-body systems with medium and strong coupling require non-perturbative approaches, in contrast to weakly coupled ones, which can be reasonably tackled by perturbation theory with a small parameter, typically associated with the interaction strength. A non-perturbative treatment can be organized by rearranging the many-body Hamiltonian in a way the perturbation theory can still be applied or, alternatively, by different criteria. In both cases, an adequate quantification of microscopic mechanisms of emergent collectivity and its thermal evolution plays an important role in solving the quantum many-body problem at different scales.
@@
It is recognized that the nuclear response theory capable of quantitatively addressing nuclear spectra must contain in-medium induced NN interaction, which is beyond the reach of the random phase approximation (RPA) Bohm and Pines (1951), or its superfluid variant, the quasiparticle RPA (QRPA) Baranger (1960); Ring and Schuck (1980). The leading effects are attributed to the minimal coupling between the single-particle and emergent collective degrees of freedom Bohr and Mottelson (1969, 1975); Broglia and Bortignon (1976); Bortignon et al. (1977); Bertsch et al. (1983); Soloviev (1992) which cause fragmentation of the (Q)RPA states or damping of collective excitations. On the formal level, the quasiparticle-vibration coupling (qPVC) can be linked to the dynamical kernels of the equations of motion (EOM) for the CFs in nuclear medium Ring and Schuck (1980); Schuck (1976); Adachi and Schuck (1989); Danielewicz and Schuck (1994); Dukelsky et al. (1998); Schuck and Tohyama (2016); Litvinova and Schuck (2019, 2020). In the regime of the NN intermediate coupling, qPVC emerges with a new order parameter associated with the qPVC vertex, which can be used to produce a hierarchy of approximations to nucleonic propagators. The leading-order qPVC kernels can be mapped to the kernels of nuclear field theory (NFT) and related to the formalism of the quasiparticle-phonon models (QPM) Bohr and Mottelson (1969, 1975); Broglia and Bortignon (1976); Bortignon et al. (1977); Bertsch et al. (1983); Tselyaev (1989); Kamerdzhiev et al. (1997); Tselyaev (2013); Soloviev (1992) in the fully microscopic EOM framework Litvinova and Schuck (2019, 2020). The latter allowed for establishing a connection between the ab initio and effective theories and extending NFT to more complex configurations.
@@
In Ref. Litvinova and Zhang (2022), the two-fermion EOM approach was advanced to the superfluid case by working out the complete formalism in the space of the Bogoliubov quasiparticles and keeping the $\times2$ matrix structure of the Hartree-Fock-Bogoliubov (HFB) basis. As in the non-superfluid phase, the modification of the bare interaction in the strongly correlated medium was derived, revealing the splitting of the interaction kernel into the static and dynamical components, in principle, calculable from the underlying bare interactions. The static kernel is associated with short-range correlations, while its dynamical counterpart generates long-range correlations originating from the retardation effects of the correlated medium, and these correlations, in turn, affect the static kernel. The major theoretical step made in Ref. Litvinova and Zhang (2022) was the unification of the particle-hole (phph) and particle-particle (pppp) channels in one EOM with the interaction kernel unifying the normal and pairing phonons and thus extending the applicability of the theory to open-shell nuclei, which constitute the major part of the nuclear landscape.
@@
The finite-temperature variants of the response theory beyond QRPA were worked out for the phph and pppp channels separately in Refs. Litvinova and Wibowo (2018); Wibowo and Litvinova (2019); Litvinova et al. (2020) and Litvinova and Schuck (2021), respectively. Applications to the electric dipole strength Litvinova and Wibowo (2018); Wibowo and Litvinova (2019), beta decay Litvinova et al. (2020), and electron capture rates Litvinova and Robin (2021) revealed the necessity of completion of the theory by developing a finite-temperature formalism including superfluidity in a unified way. This is done in the present work coherently. We derive and discuss the most general EOM and leading approximations to the interaction kernel. We benchmark the static kernel of the superfluid response by the finite-temperature QRPA Sommermann (1983); Ring et al. (1984) and then focus on the qPVC variants of the dynamical kernel, bringing them to the form ready for numerical implementations, which will be presented in a separate article.
@@ II.1.1 Zero temperature
with ${\bar{v}}_{1234}=v_{1234}-v_{1243}$ being the antisymmetrized matrix elements of the interaction between two fermions in free space. For a self-bound many-body system, $V$ in the analytical form or in the form of the two-body matrix elements should define the properties of the system, in principle, uniquely. Here, we assume only the instantaneous character of the interaction without detailing its other properties. This assumption should be reasonable enough for the low-energy fermionic systems of interest. A superfluid system of fermions is most conveniently described in the space of Bogolyubov’s quasiparticles, which mix particle and hole states via the transformation Bogolyubov (1958); Ring and Schuck (1980)
@@
$V$ $=$ $\sum\limits_{\mu\mu^{\prime}\nu\nu^{\prime}}\bigl(H^{40}_{\mu\mu^{\prime}\nu\nu^{\prime}}\alpha^{\dagger}_{\mu}\alpha^{\dagger}_{\mu^{\prime}}\alpha^{\dagger}_{\nu}\alpha^{\dagger}_{\nu^{\prime}}+\text{h.c}\bigr)$ (14)
$+$ $\sum\limits_{\mu\mu^{\prime}\nu\nu^{\prime}}\bigl(H^{31}_{\mu\mu^{\prime}\nu\nu^{\prime}}\alpha^{\dagger}_{\mu}\alpha^{\dagger}_{\mu^{\prime}}\alpha^{\dagger}_{\nu}\alpha_{\nu^{\prime}}+\text{h.c}\bigr)$
@@
$+{\tilde{F}}^{11}_{\mu\mu^{\prime}}\alpha^{\dagger}_{\mu^{\prime}}\alpha_{\mu}\Bigr)+\sum_{\mu}F^{00}_{\mu},\ \ \ \ \ \ \$
@@
$\times\left(\begin{array}[]{cccc}X^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&X^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] Y^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&Y^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] U^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&U^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\\[3.0pt] V^{n}_{\mu\mu^{\prime}}X^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}Y^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}U^{n\ast}_{\nu\nu^{\prime}}&V^{n}_{\mu\mu^{\prime}}V^{n\ast}_{\nu\nu^{\prime}}\end{array}\right)\left(\begin{array}[]{c}F^{02\ast}_{\nu\nu^{\prime}}\\[3.0pt] F^{20\ast}_{\nu\nu^{\prime}}\\[3.0pt] F^{11\ast}_{\nu\nu^{\prime}}\\[3.0pt] {\tilde{F}}^{11\ast}_{\nu\nu^{\prime}}\end{array}\right)$ (29)
@@
Thanks to the block matrix notations, the EOM for $^{\hat{R}}$ can be derived keeping the $4\times 4$ structure consistently. Namely, we consider the derivative with respect to the first time argument:
@@
${H},A](t)=e^{i{H}t}[{H},A]e^{-i{H}t}$ (44)
for an arbitrary single-time operator (product) AA. At this step, we evaluate explicitly the commutators with the one-quasiparticle part of the full Hamiltonian (13):
@@
$H,{\hat{\Psi}}_{{\mu\mu^{\prime}}}]=-{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}+[V,{\hat{\Psi}}_{{\mu\mu^{\prime}}}]$ (46)
@@
$+i\langle T[V,{\hat{\Psi}}_{\mu\mu^{\prime}}](t){\hat{\Psi}}^{\dagger}_{\nu\nu^{\prime}}(t^{\prime})\rangle,\ \ \ \ \ \ \ \ \ \$(55)
@@
${\hat{F}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t$ - $t^{\prime})[-i\overleftarrow{\partial_{t^{\prime}}}-{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}]=$ (59)
@@
where ${\hat{T}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ and ${\hat{T}}^{r}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t-t^{\prime})$ introduce the generalized static and dynamical (retarded, or time-dependent) parts of the two-fermion $^{\hat{T}}$-matrix in the quasiparticle space:
@@
$\left[i\partial_{t}\right$. - $\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}\right]{\hat{R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(t-t^{\prime})\left[-i\overleftarrow{\partial_{t^{\prime}}}-{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}\right]$ (62)
@@
$\left[\omega\right$. - $\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\mu\mu^{\prime}}\right]{\hat{R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega)\left[\omega\right.-\left.{\hat{\Sigma}}_{3}{\hat{E}}_{\nu\nu^{\prime}}\right]$
@@
${\hat{R}}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega)={\hat{R}}^{0}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega)\ \ \ \ \ \ \ \ \ \ \ \ \$
$+\frac{1}{4}\sum\limits_{\gamma\gamma^{\prime}\delta\delta^{\prime}}{\hat{R}}^{0}_{\mu\mu^{\prime}\gamma\gamma^{\prime}}(\omega){\hat{T}}_{\gamma\gamma^{\prime}\delta\delta^{\prime}}(\omega){\hat{R}}^{0}_{\delta\delta^{\prime}\nu\nu^{\prime}}(\omega)\ \ \ \ \$(65)
@@
Transformation of the $T$-matrix equation to the Dyson-Bethe-Salpeter equation (Dyson-BSE) form can be performed by introducing the interaction kernel ${\hat{K}}(\omega)={\hat{K}}^{0}+{\hat{K}}^{r}(\omega)$, irreducible with respect to ${\hat{R}}^{0}$, i.e. $^{\hat{K}}$ does not contain terms partitioned by the uncorrelated propagator ${\hat{R}}^{0}$. The interaction kernel plays the role of the self-energy for the two-point two-quasiparticle correlation function: